# Fast Direct: Query-Efficient Online Black-box Guidance for Diffusion-model Target Generation

### v4

@@ 4.1 Image Black-box Target Generation Task
Problem: Prompt Alignment. We consider the image-prompt alignment problem. While the current state-of-the-art image generative models excel at generating highly realistic images, they sometime struggle to faithfully generate images that accurately aligned with the input prompts, especially those complex prompts involving rare object combinations, object counting, or specific object positioning.
Pre-trained Model: SDXL-Lightning. We use SDXL (Podell et al., 2023) diffusion model as the backbone text-to-image model. It is able to generate 1024$\times$1024 high resolution realistic image. In our experiment, we use the distilled version, SDXL-Lightning (Lin et al., 2024), for its high sampling efficiency, which can generate image with comparable quality with just $K=8$ steps. We use the official implementation  [^2] https://huggingface.co/ByteDance/SDXL-Lightning
@@
The query to Gemini 1.5 composes of the generated images with the question like: ”Does the prompt \$prompt accurately describe the image? Rate from 1 to 5”. We state the complete query in Appendix G.
@@
Because it is a closed-source paid service, we limit the number of batch queries in our experiments, referred to as the batch queries budget. We use the Gemini 1.5 Flash model (code: gemini-1.5-flash-001) for its cost-efficiency, and set the temperature as 00 for experiments consistency and reproducibility. For conciseness, we call it simply Gemini in the following section.
@@
Experiment Procedure. We identify 12 prompts that the pre-trained model SDXL-Lighning struggles to generate, and refer these as the 12 tasks in our experiment, where the goal is to generate images that accurately aligned with the input prompts. We compare our Fast Direct algorithm against each baselines methods, and each experiment is constrained with $50$ of batch queries budget.
@@
We perform inference-time guidance using Fast Direct (Algorithm 2) to maximize the Gemini rating. We use the GP model in Eq. 2 for the pseudo-target, and set the kernel as Gaussian kernel, and we follow (Hvarfner et al., 2024) to set the lengthscale as $\lambda=\sqrt{d}$, where $d=4\times 128\times 128$ is the latents dimensionality of SDXL. We use $N=50$ iterations to utilize $50$ batch queries budget, and set the batch size as $B=32$ and step size as $\alpha=80$. We use the EularDescreteScheduler (Karras et al., 2022) sampler as suggested by the SDXL-Lightning implementation (Lin et al., 2024), and the DDIMScheduler (Song et al., 2020) (DDIM) sampler for a fair comparison with the baselines.
@@
For all baseline methods, we perform $50$ iterations, each iteration utilizes one batch query for updating. We set the batch size as $32$, and leave the rest of the hyper-parameters as default. We state more details for baselines in the Appendix D. Note that for DNO, in which each experiment trial can only produce one image, we run 16 independent experiment trials and report the average results [^3] Note that DNO spent $16\times 50=800$ batch queries for each task..
@@
Experiment Results. We present the results for the ”deer-elephant”  [^4] The complete prompt is: A yellow reindeer and a blue elephant. task in this section and defer the results for the other 11 tasks to Appendix J due to space constraints. We compare the images generated by our method with those generated by the baseline methods at each $10$ batch queries intervals in Fig. 2. We can observe that our algorithm successfully satisfies the visual objectives for both samplers within the 50 batch queries budget, while the baseline methods fall short within the same budget. Similar phenomenon are observed for the other 11 tasks in Appendix J. We grant baseline methods with extra batch queries budget, and we observed that DDPO and DPOK almost achieves the visual objectives, whereas D3PO and DNO remain unsuccessful in this particular task.
@@
We collect the 32 randomly generated images using our algorithm (by utilizing $50$ batch queries budget) and present in Fig. 3. We can observe that most images are well satisfies the visual objective. Similar phenomenon are observed on the other 11 tasks as shown in Appendix J.
@@
Figure 2: The generated images over each number of batch queries on the prompt ”deer-eleplant” 4, extra batch queries budgets (until 500) are given to the baselines methods for demonstration.
@@
Figure 3: The 32 randomly generated image for the prompt ”deer-eleplant” 4 guided by Fast Direct (w/ EDM) by utilizing 50 batch queries budget.
@@
For the quantitative results, we report the average objective values over each batch query across the 12 tasks in Fig. 4. For DNO, which only produces one image each experiment, we average the results from 16 independent experiments. We observe that the Fast Direct achieves nearly full score within the 50 batch queries budget in all tasks, which is significant higher than the baseline methods. Moreover, we observe that Fast Direct achieve similar score for both samplers, suggesting it is invariant to sampler. We further perform ablation study in Appendix B, and show that Fast Direct is insensitive to the hyper-parameters: step size $\alpha$ and batch size $B$.
We further employ tasks 1 to 3 and assigns extra batch queries budget of $200$ to the baseline methods. We report the accumulated objective values, which represents the best objective values achieved up to each number of batch queries, in Appendix Fig. 11. We observe that our algorithm, utilizing only 50 batch queries budget, consistently outperforms all baseline methods even when they are assigned with expanded batch queries budget of 200.
To quantify this advantage, in Table 1, we identify the minimum number of batch queries budget (denotes as $N^{*}$) our algorithm requires to surpass each baseline methods with their $200$ batch query budget. We further calculate ours batch-query-efficiency gain compared to baselines as $\frac{200}{N^{*}}$ and report (in the bracket) in Table 1. We can observe that our algorithm is at least 667% and up to 1000% more batch-queries-efficient than the baseline methods.
Table 1: The minimum number of batch queries budget (denotes as $N^{*}$) for Fast Direct (ours) to outperform each baselines (when each baseline is granted with $200$ batch queries budget) on images and molecules tasks. The values in bracket show the batch-query-efficiency gain ($\frac{200}{N^{*}}$) of Fast Direct (ours) over the baselines.
@@ 4.2 Molecule Black-box Target Generation Task
Problem: Drug Discovery. We consider the drug discovery problem. One of the key problem is to find drug molecules that has a strong binding affinity with the target protein receptor. The binding affinity quantify the strength of the interaction between two bio-molecules, so a stronger binding affinity indicates a higher drug efficacy.
@@
A traditional approach to this problem requires human-design molecules, which are highly time-consuming. Recent advancements in AI poses a great potential to speedup this process by generating molecules using advanced generative model such as diffusion model. Although the current pre-trained molecules models, such as TargetDiff, are capable of generating relatively realistic molecules, it lacks the ability of generating molecules with user desired high binding affinity.
@@
Pre-trained Model: TargetDiff. We use TargetDiff (Guan et al., 2023) as the backbone molecules generative model. It is pre-trained on the CrossDocked2020 dataset (Francoeur et al., 2020), and can generate realistic 3D molecular structures conditioned on the given protein receptor. For sampling efficiency, we use the DDIMScheduler sampler with $K=200$ sampling steps as we find that it can generate comparable results. During the generation process, we fix the atoms type according to the dataset reference and only allow the atoms position to be varied. We use the official implementation  [^5] https://github.com/guanjq/targetdiff in our experiment.
@@
Objective Function: Molecules Binding Affinity. Measurement of the binding affinity requires real-world experiment which is highly expensive (David et al., 2020). In our experiment, we leverage the Vina score (kcal/mol) calculated by the AutoDock Vina simulation software (Eberhardt et al., 2021) to estimate the binding affinity. A lower (more negative) Vina score indicates a stronger estimated binding affinity. Therefore, our objective is to minimize the Vina score.
@@
Experiment Procedure. We conduct experiments on 6 tasks, where the goal of each task is to optimize the Vina score on the protein receptors of ID from 1 to 6, respectively. Similar to images tasks, each algorithm is constrained with 50 of batch queries budget.
@@
We perform inference-time guidance using our Algorithm 2 to minimize the Vina score. We use the historical optimal in Eq. 4 for the pseudo-target. We use $N=50$ iterations to utilize $50$ batch queries budget, and set the batch size as $B=32$, and the step size as $\alpha=10^{-2}$. We then conduct experiments on the baseline methods for $50$ fine-tuning iteration. For DDPO and DPOK, the batch size is set as the same $32$; while D3PO requires binary rewards, so we double the batch size as $64$  [^6] However, unlike image tasks where Gemini can evaluate rewards for two images in a single query, the Vina simulation lacks this capability, requiring 64 individual queries for a batch of size 64.. We keep the hyper-parameters as default for all baseline methods.
@@
Experiment Results. We report the average Vina score (recall, lower is better) for each batch queries across 6 tasks in Fig. 10. We observe that our algorithm consistently achieves much lower Vina score compared to the baseline methods across all tasks. Similar to the images task, we grant extra batch queries budget of 200 to the baselines for task 1 and 2, and report the accumulated objective values in Appendix Fig 12, and list the batch-query-efficiency gain in Table 1. We observe that our algorithm is at least $1$% and up to $4$% more batch-query-efficient compared to the baseline methods.
@@ 5 Conclusion
In this work, we proposed Fast Direct for diffusion model target generation, which effectively addresses the challenges of limited batch query budgets and black-box objective functions, demonstrating its potential for various applications, including image generation and drug discovery. Fast Direct is highly practical, as it is easy to implement, supports any type of SDE solver, and has only one hyper-parameter to tune (step size $\alpha$). Our algorithm is based on the surprising empirical observation that the universal update direction (i.e., $\bm{x}^{*}-\bm{x}_{K}$ in Algorithm 1) can efficiently guide the diffusion trajectory toward the target, even when the target $\bm{x}^{*}$ is extremely noisy. This phenomenon suggests an intriguing robustness in the diffusion inference process. Future research could investigation for its underlying theoretical principles.
@@ Acknowledgments
This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative, the Center for Frontier Artificial Intelligence Research Center, Institute of High Performance Computing, A*Star, and the College of Computing and Data Science at Nanyang Technological University. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority. Yueming Lyu is supported by Career Development Fund (CDF) of the Agency for Science, Technology and Research (A*STAR) (No: C233312007).
@@ Appendix A More Experiments: Compressibility, Incompressibility, and Aesthetic Quality
Objective Function. For compressibility, we aimed to minimize the compressed JPEG size (MB) of the generated images. For incompressibility, it’s simply the inverse. For aesthetic quality, the aesthetic score is evaluated by the pre-trained LAION aesthetics predictor (schuhmann2022laion), which is trained on human rating of the images aesthetic quality, and we aimed to maximize the aesthetic score.
Experiment Procedure. For each task, we uniformly sample from the 45 common animals (as proposed by DDPO (Black et al., 2023)) as input prompts and optimize the objective using Fast Direct in Algorithm 2. The prompt for each instance within a batch is sampled randomly and independently. We set $N=50$ batch queries budget for compressibility and incompressibility, and $N=100$ for aesthetic quality. We use the same pre-trained model, hyper-parameters, and GP settings as in Section 4.1. The 45 common animals prompts used in the experiment is as follows:
@@
For DNO, where each experimental trial generates only a single image and the objective score is highly dependent on the specific prompt. To ensure a more accurate evaluation, we run $45$ independent experiment trials using each prompt and report the average results. Consequently, DNO requires 45×45$\times$45 × more batch queries per task compared to other methods.
@@
Figure 5: Left column: The average objective score of the generated images over each number of batch queries on the 3 black-box optimization tasks, the images are generated using the 45 common animals that used in DDPO (Tang et al., 2024) as the input prompts. Right column: The objective score of the images generated by unseen prompts, it demonstrates the generalization capability. Note that DNO is not applicable to this task.
Generalization Ability. We follow DDPO to evaluate generalization ability. For Fast Direct, we freeze the learned GP model to generate 16 unseen images using distinct unseen animal prompts. Specifically, in this phase, Lines 17 and 18 are removed from Algorithm 2, and our algorithm does not access the objective function. For DDPO, DPOK, and D3PO, the fine-tuned models are frozen to generate images with the 16 unseen animal prompts. DNO is not applicable to this experiment because DNO needs to perform optimization for each image without generalization. For the unseen prompts, since DDPO didn’t publish the unseen prompts, we created our own as follows:
@@
In Fig.6, we present the images generated using the unseen prompt ”hippo” by each algorithm across the three tasks. The remaining 15 prompts for the three tasks are provided in the supplementary material. A similar phenomenon is observed for all 15 other unseen prompts across the three tasks. For quantitative evaluation, we report the objective values for the unseen prompts in the right column of Fig.5. We can observe that Fast Direct with GP model has generalization ability to unseen prompts.
@@
Figure 6: The generated images using the unseen prompt ”hippo” in aesthetic quality task, extra batch queries budgets (until 500) are given to the baselines methods for demonstration.
@@
Figure 7: The generated images using the 16 unseen prompts. Top left is the images generated by pre-trained model, the rest are the Fast Direct generated images in the corresponding tasks.

### v5

@@ 4.1 Image Black-box Target Generation Task
Problem: Prompt Alignment. We consider the image-prompt alignment problem. While the current state-of-the-art image generative models excel at generating highly realistic images, they sometimes struggle to faithfully generate images that are accurately aligned with the input prompts, especially those complex prompts involving rare object combinations, object counting, or specific object positioning.
Pre-trained Model: SDXL-Lightning. We use the SDXL (Podell et al., 2023) diffusion model as the backbone text-to-image model. It can generate a 1024$\times$1024 high-resolution realistic image. In our experiment, we use the distilled version, SDXL-Lightning (Lin et al., 2024), for its high sampling efficiency, which can generate images with comparable quality with just $K=8$ steps. We use the official implementation  [^2] https://huggingface.co/ByteDance/SDXL-Lightning
@@
The query to Gemini 1.5 consists of the generated images with the question like: ”Does the prompt \$prompt accurately describe the image? Rate from 1 to 5”. We state the complete query in Appendix G.
@@
Because it is a closed-source paid service, we limit the number of batch queries in our experiments, referred to as the batch query budget. We use the Gemini 1.5 Flash model (code: gemini-1.5-flash-001) for its cost-efficiency, and set the temperature to 00 for experiment consistency and reproducibility. For conciseness, we call it simply Gemini in the following section.
@@
Experiment Procedure. We identify 12 prompts that the pre-trained model SDXL-Lighning struggles to generate, and refer to these as the 12 tasks in our experiment, where the goal is to generate images that are accurately aligned with the input prompts. We compare our Fast Direct algorithm against each baseline method, and each experiment is constrained with a $50$ batch query budget.
@@
We perform inference-time guidance using Fast Direct (Algorithm 2) to maximize the Gemini rating. We use the GP model in Eq. 2 for the pseudo-target, and set the kernel as a Gaussian kernel, and we follow (Hvarfner et al., 2024) to set the lengthscale as $\lambda=\sqrt{d}$, where $d=4\times 128\times 128$ is the latent dimensionality of SDXL. We use $N=50$ iterations to utilize the $50$ batch query budget, and set the batch size as $B=32$ and step size as $\alpha=80$. We use the EularDescreteScheduler (Karras et al., 2022) sampler as suggested by the SDXL-Lightning implementation (Lin et al., 2024), and the DDIMScheduler (Song et al., 2020) (DDIM) sampler for a fair comparison with the baselines.
@@
For all baseline methods, we perform $50$ iterations, each iteration utilizes one batch query for updating. We set the batch size to $32$, and leave the rest of the hyperparameters as default. We state more details for baselines in the Appendix D. Note that for DNO, in which each experiment trial can only produce one image, we run 16 independent experiment trials and report the average results [^3] Note that DNO spent $16\times 50=800$ batch queries for each task..
@@
Experiment Results. We present the results for the ”deer-elephant”  [^4] The complete prompt is: A yellow reindeer and a blue elephant. task in this section and defer the results for the other 11 tasks to Appendix J due to space constraints. We compare the images generated by our method with those generated by the baseline methods at each $10$ batch query intervals in Fig. 2. We can observe that our algorithm successfully satisfies the visual objectives for both samplers within the 50 batch query budget, while the baseline methods fall short within the same budget. Similar phenomena are observed for the other 11 tasks in Appendix J. We grant baseline methods with extra batch query budget, and we observed that DDPO and DPOK almost achieve the visual objectives, whereas D3PO and DNO remain unsuccessful in this particular task.
@@
We collect the 32 randomly generated images using our algorithm (by utilizing a $50$ batch query budget) and present them in Fig. 3. We can observe that most images well satisfy the visual objective. Similar phenomena are observed on the other 11 tasks as shown in Appendix J.
@@
Figure 2: The generated images over each number of batch queries on the prompt ”deer-eleplant” 4, extra batch query budget (until 500) is given to the baseline methods for demonstration.
@@
Figure 3: The 32 randomly generated images for the prompt ”deer-eleplant” 4 guided by Fast Direct (w/ EDM) by utilizing 50 batch query budget.
@@
For the quantitative results, we report the average objective values over each batch query across the 12 tasks in Fig. 4. For DNO, which only produces one image each experiment, we average the results from 16 independent experiments. We observe that the Fast Direct achieves nearly a full score within the 50 batch query budget in all tasks, which is significantly higher than the baseline methods. Moreover, we observe that Fast Direct achieve a similar score for both samplers, suggesting it is invariant to the sampler. We further perform an ablation study in Appendix B, and show that Fast Direct is insensitive to the hyper-parameters: step size $\alpha$ and batch size $B$.
We further employ tasks 1 to 3 and assign an extra batch query budget of $200$ to the baseline methods. We report the accumulated objective values, which represent the best objective values achieved up to each number of batch queries, in Appendix Fig. 11. We observe that our algorithm, utilizing only 50 batch query budget, consistently outperforms all baseline methods even when they are assigned with expanded batch query budget of 200.
To quantify this advantage, in Table 1, we identify the minimum number of batch query budget (denoted as $N^{*}$) our algorithm requires to surpass each baseline method with their $200$ batch query budget. We further calculate ours batch-query-efficiency gain compared to baselines as $\frac{200}{N^{*}}$ and report (in the bracket) in Table 1. We can observe that our algorithm is at least 667% and up to 1000% more batch-query-efficient than the baseline methods.
Table 1: The minimum number of batch query budget (denoted as $N^{*}$) for Fast Direct (ours) to outperform each baseline (when each baseline is granted with $200$ batch query budget) on images and molecules tasks. The values in brackets show the batch-query-efficiency gain ($\frac{200}{N^{*}}$) of Fast Direct (ours) over the baselines.
@@ 4.2 Molecule Black-box Target Generation Task
Problem: Drug Discovery. We consider the drug discovery problem. One of the key problems is to find drug molecules that have a strong binding affinity with the target protein receptor. The binding affinity quantifies the strength of the interaction between two biomolecules, so a stronger binding affinity indicates a higher drug efficacy.
@@
A traditional approach to this problem requires human-designed molecules, which are highly time-consuming. Recent advancements in AI pose a great potential to speed up this process by generating molecules using advanced generative models such as diffusion models. Although the current pre-trained molecule models, such as TargetDiff, are capable of generating relatively realistic molecules, they cannot generate molecules with the user’s desired high binding affinity.
@@
Pre-trained Model: TargetDiff. We use TargetDiff (Guan et al., 2023) as the backbone molecules generative model. It is pre-trained on the CrossDocked2020 dataset (Francoeur et al., 2020), and can generate realistic 3D molecular structures conditioned on the given protein receptor. For sampling efficiency, we use the DDIMScheduler sampler with $K=200$ sampling steps as we find that it can generate comparable results. During the generation process, we fix the atom type according to the dataset reference and only allow the atom position to be varied. We use the official implementation  [^5] https://github.com/guanjq/targetdiff in our experiment.
@@
Objective Function: Molecules Binding Affinity. Measurement of the binding affinity requires a real-world experiment, which is highly expensive (David et al., 2020). In our experiment, we leverage the Vina score (kcal/mol) calculated by the AutoDock Vina simulation software (Eberhardt et al., 2021) to estimate the binding affinity. A lower (more negative) Vina score indicates a stronger estimated binding affinity. Therefore, our objective is to minimize the Vina score.
@@
Experiment Procedure. We conduct experiments on 6 tasks, where the goal of each task is to optimize the Vina score on the protein receptors of ID from 1 to 6, respectively. Similar to image tasks, each algorithm is constrained to 50 batch query budget.
@@
We perform inference-time guidance using our Algorithm 2 to minimize the Vina score. We use the historical optimal in Eq. 4 for the pseudo-target. We use $N=50$ iterations to utilize the $50$ batch query budget, and set the batch size as $B=32$, and the step size as $\alpha=10^{-2}$. We then conduct experiments on the baseline methods for $50$ fine-tuning iterations. For DDPO and DPOK, the batch size is set as the same $32$; while D3PO requires binary rewards, so we double the batch size as $64$  [^6] However, unlike image tasks where Gemini can evaluate rewards for two images in a single query, the Vina simulation lacks this capability, requiring 64 individual queries for a batch of size 64.. We keep the hyperparameters as default for all baseline methods.
@@
Experiment Results. We report the average Vina score (recall, lower is better) for each batch of queries across 6 tasks in Fig. 10. We observe that our algorithm consistently achieves much lower Vina score compared to the baseline methods across all tasks. Similar to the images task, we grant extra batch query budget of 200 to the baselines for tasks 1 and 2, and report the accumulated objective values in Appendix Fig 12, and list the batch-query-efficiency gain in Table 1. We observe that our algorithm is at least $1$% and up to $4$% more batch-query-efficient compared to the baseline methods.
@@ 5 Conclusion
In this work, we proposed Fast Direct for diffusion model target generation, which effectively addresses the challenges of limited batch query budgets and black-box objective functions, demonstrating its potential for various applications, including image generation and drug discovery. Fast Direct is highly practical, as it is easy to implement, supports any type of SDE solver, and has only one hyper-parameter to tune (step size $\alpha$). Our algorithm is based on the surprising empirical observation that the universal update direction (i.e., $\bm{x}^{*}-\bm{x}_{K}$ in Algorithm 1) can efficiently guide the diffusion trajectory toward the target, even when the target $\bm{x}^{*}$ is extremely noisy. This phenomenon suggests an intriguing robustness in the diffusion inference process. Future research could investigate its underlying theoretical principles.
@@ Acknowledgments
This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative, the Center for Frontier Artificial Intelligence Research Center, Institute of High Performance Computing, A*Star, and the College of Computing and Data Science at Nanyang Technological University. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority. Yueming Lyu is supported by Career Development Fund (CDF) of the Agency for Science, Technology and Research (A*STAR) (No: C243512014 and No: C233312007).
@@ Appendix A More Experiments: Compressibility, Incompressibility, and Aesthetic Quality
Objective Function. For compressibility, we aimed to minimize the compressed JPEG size (MB) of the generated images. For incompressibility, it’s simply the inverse. For aesthetic quality, the aesthetic score is evaluated by the pre-trained LAION aesthetics predictor (schuhmann2022laion), which is trained on human ratings of the images’ aesthetic quality, and we aimed to maximize the aesthetic score.
Experiment Procedure. For each task, we uniformly sample from the 45 common animals (as proposed by DDPO (Black et al., 2023)) as input prompts and optimize the objective using Fast Direct in Algorithm 2. The prompt for each instance within a batch is sampled randomly and independently. We set $N=50$ batch query budget for compressibility and incompressibility, and $N=100$ for aesthetic quality. We use the same pre-trained model, hyper-parameters, and GP settings as in Section 4.1. The 45 common animal prompts used in the experiment are as follows:
@@
For DNO, each experimental trial generates only a single image, and the objective score is highly dependent on the specific prompt. To ensure a more accurate evaluation, we run $45$ independent experiment trials using each prompt and report the average results. Consequently, DNO requires 45×45$\times$45 × more batch queries per task compared to other methods.
@@
Figure 5: Left column: The average objective score of the generated images over each number of batch queries on the 3 black-box optimization tasks, the images are generated using the 45 common animals that were used in DDPO (Tang et al., 2024) as the input prompts. Right column: The objective score of the images generated by unseen prompts, which demonstrates the generalization capability. Note that DNO is not applicable to this task.
Generalization Ability. We follow DDPO to evaluate generalization ability. For Fast Direct, we freeze the learned GP model to generate 16 unseen images using distinct unseen animal prompts. Specifically, in this phase, Lines 17 and 18 are removed from Algorithm 2, and our algorithm does not access the objective function. For DDPO, DPOK, and D3PO, the fine-tuned models are frozen to generate images with the 16 unseen animal prompts. DNO is not applicable to the generalization experiment, as it performs per-image optimization and lacks generalization ability. For the unseen prompts, since DDPO didn’t publish the unseen prompts, we created our own as follows:
@@
In Fig.6, we present the images generated using the unseen prompt ”hippo” by each algorithm across the three tasks. The remaining 15 prompts for the three tasks are provided in the supplementary material. A similar phenomenon is observed for all 15 other unseen prompts across the three tasks. For quantitative evaluation, we report the objective values for the unseen prompts in the right column of Fig.5. We can observe that Fast Direct with the GP model has generalization ability to unseen prompts.
@@
Figure 6: The generated images using the unseen prompt ”hippo” in the aesthetic quality task, extra batch query budgets (until 500) are given to the baseline methods for demonstration.
@@
Figure 7: The generated images using the 16 unseen prompts. Top left is the image generated by the pre-trained model, the rest are the Fast Direct generated images in the corresponding tasks.